OpenStudy (anonymous):

Write the equation of the quadratic function with roots -10 and -8 and a vertex at (-9, -3).

OpenStudy (anonymous):

Hint: Vertex form of a quadratic equation \[f(x) = a(x-h)^2+k\] where your vertex is (-9,-3)

OpenStudy (anonymous):

Sorry I meant to say where your vertex is (h,k) but you get the point :P

OpenStudy (campbell_st):

or you can use if x = -10 is a root, then (x + 10) is a factor x = -8 is a root, then (x + 8) is a factor the the quadratic is y = a(x +10)(x + 8) then substitute the vertex to find the value of a

OpenStudy (anonymous):

@Coolguy123